OpenStudy (anonymous):

How can you tell when a quadratic equation has two identical, rational solutions?

OpenStudy (bahrom7893):

Discriminant = 0

OpenStudy (bahrom7893):

or b^2 - 4ac = 0 (for an equation in the form of: ax^2+bx+c = 0)

OpenStudy (anonymous):

thank you

hero (hero):

bahrom, it would have been nice to see an example

OpenStudy (anonymous):

if \[\Delta=b^2-4ac=0\] then you have just one solution, because roots are the same..

OpenStudy (bahrom7893):

hero, you can provide one.

OpenStudy (anonymous):

if \[\Delta >0\] then you have two real solution

hero (hero):

Why do I have to provide an example for a general solution that I didn't provide? I'm asking you to prove that what you say is true.